OpenStudy (anonymous):

The formula below gives I, the interest earned by depositing p dollars in a bank account at annual interest rate r for one year. I = pr If a deposit earns $75 of interest, which equation gives the principal p as a function of the interest rate r? Can Someone please help

OpenStudy (solomonzelman):

it sounds like your question has options to it (we don't need the options to solve it, but if there are options, please post them)

OpenStudy (solomonzelman):

if you need a function of p, in terms of r, when p=75, then: divide both sides by r, and mlup in 75 for p

OpenStudy (solomonzelman):

*plug, not mlup

OpenStudy (anonymous):

here are the answers... A. p = 75r B. p = r/75 C. p = 75 – r D. p = 75/r

OpenStudy (anonymous):

@SolomonZelman

OpenStudy (anonymous):

try the triangle trick. it always helps me.

OpenStudy (anonymous):